3. Two pulley 12 in diameter and 20 in diameter is connected by a crossed belt with center distance of 45 in. Find the belt length needed to connect them.


1
Expert's answer
2022-03-17T14:56:04-0400

Length of the belt :

"L = \pi(r_1+r_2)+2x+\frac{(r_1+r_2)^2}{x}"

"L = \pi(10+6)+2\times45+\frac{(10+6)^2}{45} =146" in